HETAV PATEL

Grade 11,

what are the topics in which we can score good marks

what are the topics in which we can score good marks

Grade:12th pass

2 Answers

Deepak Kumar Shringi
5 years ago
Ganesh
2 years ago
The question is simple angular acceleration = Torque/I , Torque  = 2* 3 N-m and mass moment of inertia  = 1*(3^2+4^2) =25
hence answer is 2/25 = 0.88rad/s^2
contact me for clarifications at: